OpenStudy (dmezzullo):

Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (dmezzullo):

@mikala1 @jazy

OpenStudy (anonymous):

PEMDAS 2(x + 6) − 10 = 12 + 4(x − 1) first simplify 2x+12-10= 12+4x-4 add or subtract like terms 2x+2= 8+4x subtract 2 both sides -2 -2 ________________________________________ 2x= 6+4x subtract 4x on both sides -4x -4x ________________ -2x=6 divide -2 on both sides -2x/-2 = 6/-2 x= ?

OpenStudy (anonymous):

lol yea... she beat me to it...

OpenStudy (anonymous):

Aussy123 is correct.

OpenStudy (dmezzullo):

-3?

OpenStudy (anonymous):

YEP!! GJ

OpenStudy (dmezzullo):

tyvm medals for me? anyone

OpenStudy (anonymous):

I think I gave you one

OpenStudy (dmezzullo):

i know u did anyone else?